# Machine-generated text detection prevents language model collapse

### v1

@@ 4 The impact of decoding strategies on model collapse
Figure 1: Perplexity and accuracy of GPT-2 (top) and SmolLM2 (bottom) over generations of fully synthetic recursive training.
Model Decoding Diversity $\uparrow$ S-BLEU $\downarrow$ MAUVE $\uparrow$ Readability $\uparrow$
GPT-2 greedy $0.96$/$0.70$ $61.02$/$67.13$ $0.99$/$1.00$ $60.47$/$8.25$
beam search $16.78$/$10.86$ $61.54$/$67.60$ $0.91$/$1.29$ $60.97$/$17.57$
sampling $94.88$/$99.82$ $24.12$/$6.76$ $90.16$/$7.18$ $40.62$/$-10.14$
temperature $87.76$/$25.10$ $33.45$/$54.56$ $94.15$/$22.69$ $46.80$/$\mathbf{36.79}$
top-$k$ $84.57$/$70.20$ $38.81$/$\mathbf{42.14}$ $95.21$/$\mathbf{70.01}$ $51.19$/$34.32$
nucleus $92.26$/$\mathbf{86.73}$ $28.24$/$26.86$ $90.96$/$57.41$ $43.96$/$21.31$
SmolLM2 greedy $6.99$/$3.02$ $61.90$/$50.23$ $4.29$/$1.12$ $62.33$/$\mathbf{46.72}$
beam search $6.75$/$2.88$ $61.47$/$\mathbf{44.47}$ $3.86$/$1.09$ $62.38$/$44.47$
sampling $90.83$/$\mathbf{91.62}$ $37.69$/$33.33$ $91.28$/$\mathbf{83.82}$ $47.49$/$44.33$
temperature $82.96$/$24.37$ $46.21$/$61.64$ $91.48$/$21.66$ $52.78$/$64.59$
top-$k$ $82.57$/$57.35$ $48.62$/$61.18$ $95.48$/$68.41$ $55.45$/$62.82$
nucleus $87.38$/$47.36$ $42.43$/$60.85$ $94.40$/$67.37$ $51.14$/$63.71$
Human $88.79$ $42.89$ $100$ $50.34$
Table 1: Impact of decoding strategies on the generation quality of GPT-2 and SmolLM2 from generation 00 to $9$ of fully synthetic recursive training. For each model we highlight in bold the closest score to the human reference for generation $9$ ($\uparrow$ / $\downarrow$: higher / lower is better).
@@
We evaluate model collapse on the open-ended text generation task by finetuning LLMs on the Wiki- Text-2 dataset Merity et al. (2016) with GPT-2 and Hugging Face’s recently released SmolLM2 Allal et al. (2025). The Wikipedia articles are segmented into chunks of $512$ tokens where the first $256$ are used as the context ($\mathbf{x}$), and the remaining $256$ as the continuation ($\hat{\mathbf{x}}$). We conduct full finetuning for 1 epoch and to avoid information leakage between generation and training, define cross-entropy loss only on the generated sequence of each sample Dohmatob et al. (2024b). Additional details about the experimental setup can be found in Appendix B.1.
As described in section 3.2, we recursively train language models on the previous generation’s output. We evaluate a range of decoding strategies to assess the effect on model collapse, including greedy decoding ($\tau=0$), $5$-way beam search, pure sampling ($\tau=1$), temperature ($\tau=0.9$), top-$k$ (with $k=50$), and nucleus sampling ($\eta=0.95$). The hyperparameter settings for these decoding strategies were based on recommendations from prior work Holtzman et al. (2020); Shumailov et al. (2024); Garces Arias et al. (2025).
In Figure 1 we present the perplexity and evaluation accuracy on the WikiText test set for every model generation. Additionally, we obtain scores for the qualitative metrics using the outputs generated by the model (i.e. $\{\hat{\mathbf{x}}\}_{s=1}^{n}$ of $\mathcal{D}_{\text{S}}^{i}$ in Algorithm 1). We present the qualitative metrics at generation 0 and 9 in Table 1.
Deterministic decoding results in the most significant model collapse. While stochastic sampling appear to follow a linear relationship in model degradation across generations, the rate of collapse is accelerated with greedy decoding and beam search before plateauing in later generations. The data quality produced by the models provides some intuition for this as the starting metrics are significantly lower than stochastic methods with diversity and MAUVE scores less than $10\%$ for SmolLM2, evidence of repetitive and non-fluent language. Due to the poor data quality and exacerbated model collapse we choose not to carry out our subsequent experiments with deterministic strategies.
@@
Pure sampling affected the models in opposite ways. In the case of GPT-2, sampling directly from the probability distribution leads to diverse and fluent text in generation 00 but the effect of training on this data is the highest perplexity among all stochastic methods and generated text that does not resemble anything similar to the human reference texts with a final MAUVE of $7.18$. With SmolLM2, on the other hand, pure sampling led to the smallest drop in model performance and the highest quality data on average across all metrics.
@@
Recursive training with temperature leads to the most repetitive text of all stochastic methods with a ${}70\%$ decrease in diversity for both models and the furthest distribution semantically from the human reference for SmolLM2, showing significant model collapse.
Performance with top-$k$ was consistent across models, with the smallest drop in diversity and Self-BLEU, the closest resulting text to the human reference and a smaller drop in test performance compared to Nucleus sampling.
@@
In our subsequent experiments on preventing model collapse, we seek to validate that our method can work in the most extreme scenario while also facilitating direct comparison between models. For this reason, we evaluate the models using the worst performing stochastic decoding method (pure sampling for GPT-2 and temperature sampling for SmolLM2) and top-$k$ decoding due to its consistent performance across models that will allow for direct comparison.
5 Machine-generated text detector
To prevent model collapse in the setting where the origin of each training sample is unknown, we train a detector to estimate the likelihood of the text being machine-generated. In the following section, we define how we trained our model, followed by our method of weighted sampling using the detector outputs as importance scores.
Machine-generated text detection methods can be divided into neural-based and metric-based approaches. Metric-based methods use statistical features, often extracted from surrogate LLMs, to detect machine-generated text, whereas neural-based methods are implemented by fine-tuning a small pre-trained language model with a binary classification head. We decided on a neural-based approach due to their state-of-the-art performance on common benchmarks Wang et al. (2024); Li et al. (2024) and the ability to use the model output logits as a form of uncertainty.
@@ 5 Machine-generated text detector
ModernBERT 98.61 94.79 94.80 94.27 86.07 86.01
DeBERTaV3 97.10 95.43 95.42 81.68 81.18 80.96
RoBERTa 98.15 94.03 94.04 84.60 80.64 80.36
Table 2: Machine-generated text detection performance on MAGE. The F1-score and accuracy (Acc.) are calculated as macro-averages.
Figure 2: GPT2 (top) and SmolLM2 (bottom) under partially synthetic recursive training ($\alpha=1$) for 10 generations. The baseline is equivalent to training on all the data in the pool and the Oracle performance represents a perfect AI text detector that filters all synthetic samples.
@@
[…]
@@
Table 3: Comparison of data quality when using our mitigation strategy versus the baseline for GPT-2 and SmolLM2. Results are shown for top-$k$ decoding and pure sampling/temperature with different values for $\alpha$, $\beta$, and $\gamma$. The proportion of human data in the recursive training is given by $\alpha$. $\uparrow$ means higher is better and $\downarrow$ means lower is better. Positive improvements are shown in blue and negative results in red.
5.1 Model
Our detector is based on an encoder-only language model with a sequence classification head that maps the CLS token representation to logits, $\mathbf{z}_{i}$, which are converted to probabilities using a sigmoid function. As LLM training is considerably resource intensive, any data filtering or sampling methods must be able to efficiently process large quantities of data with minimal computational overhead Wenzek et al. (2020). Based on these requirements we evaluated the base variants of three model architectures with under 200 million parameters: RoBERTa Goyal et al. (2020) and DeBERTav3 He et al. (2023) due to their state-of-the-art performance in machine-generated text detection benchmarks and ModernBERT Warner et al. (2024), a very recent model that has achieved superior performance on a range of Natural Language Inference and Information Retrieval benchmarks. The added advantage of ModernBERT is the large context window (8192 tokens) and superior computational speed and memory efficiency Warner et al. (2024). More details on the models can be found in Appendix A.
Despite their strong performance, as with all modern neural networks, the confidence estimates are poorly calibrated, often resulting in overconfident predictions  Guo et al. (2017). To mitigate overconfidence in the predictions, we applied label smoothing to the target labels during training. Additionally, we implemented temperature scaling to further calibrate the model’s predictions. Given the logit vector $\mathbf{z}_{i}$, the new confidence prediction is $\hat{q}_{i}=\sigma_{\text{SM}}(\mathbf{z}_{i}/T)$ where $T$ is the temperature parameter optimised using negative log-likelihood loss on the validation set.
5.2 Training
Each model was trained and evaluated on the MAGE dataset Li et al. (2024), a popular AI text detection benchmark dataset based on 10 Natural Language Processing datasets which have been used to generate text from 27 LLMs. This dataset has already been split into train, validation and test splits in the ratio 80/10/10. We also test on their challenging out-of-distribution test set containing human-curated text from four unseen domains and machine-generated samples by an unseen model (GPT-4). Each model was finetuned for 5 epochs using a binary cross-entropy loss with early stopping. More details on the model training hyperparameters and the dataset can be found in Appendix A. The results of our evaluation can be seen in Table 2 where ModernBERT achieved the best classification performance on both test sets with an Area Under the Receiver Operating Characteristic Curve (AUC) of $98.61$ and $94.27$ respectively.
6 Preventing Model Collapse
So far we have carried out recursive training in the most extreme scenario where models are trained exclusively on the outputs of the previous generation and no human data is added during training. We now turn our focus to the partially synthetic setting, a more realistic scenario where human data make up a portion of the training dataset and the synthetic data is mixed across generations. In this setting, we explore model collapse with the assumption that the data distribution is unknown. The training dataset for generation $i$, $\mathcal{D}^{i}$, is defined in equation 3 with the three different scenarios depending on the mixing coefficients $\alpha$, $\beta$, and $\gamma\!\in\![0,1]$:
$\begin{split}\mathcal{D}^{i}\sim&\,\,{\text{sample}_{i\geq 1}}\,\,\mathcal{D}_{\text{H}},\alpha\\ &\,\,{\text{sample}_{i\geq 1}}\,\,\mathcal{D}_{\text{S}}^{i},\beta\\ &\,\,{\text{sample}_{i\geq 2}}\,\,\left\{\mathcal{D}_{\text{S}}^{i-1},\dots,\mathcal{D}_{\text{S}}^{1}\right\},\frac{\gamma}{(i-1)}\,.\end{split}$ (3)
  1. 1.
    Fully synthetic $\left(\alpha=0,\beta=1,\gamma=0\right)$: training data consists entirely of synthetic samples from the previous model generation.
  2. 2.
    Partially synthetic $\left(\alpha>0,\beta=1,\gamma=0\right)$: the same proportion of human data is added to the training data at every iteration.
  3. 3.
    Partially synthetic with cross-generational data $\left(\alpha>0,\beta>0,\gamma>0\right)$: as above but the synthetic data is accumulated across generations.
We evaluate our method in the partially synthetic setting and vary the mixing coefficients $\alpha,\beta,\gamma$ to affect the distribution of human and machine-generated data as well as the proportion of cross-generational data in the training set. We denote the sigmoid output from the detector for a sample $x_{n}$ as $q\left(x_{n}\right)$. Given a dataset $\mathcal{D}$, consisting of an unknown distribution of human and synthetic samples, and the predictions from the detector for each sample $q\left(x_{n}\right)\forall\mathcal{D}$, the goal of our method is to sample a new dataset which when used to finetune model $p_{\theta}$ does not result in model collapse. A model has collapsed if the addition of synthetic samples in the training data leads to degraded model performance compared to training exclusively on the human samples. We evaluate model performance and data quality as described in Sections 3.3 and 3.4.
6.1 Method
Our method is based on weighted sampling with replacement where the weights are extracted from the AI text detector model probabilities. As the detector has been trained on an unbalanced dataset ($71\%$ machine-generated), the predictions are biased towards attributing text as machine-generated, reflected in the optimal classification threshold of $0.8674$. To shift the probabilities in favour of predicting human samples, we apply a bias term (calculated based on the confidence threshold) to the probabilities, $b$, followed by normalising the weights so they sum to 1, as in
$w_{i}=\frac{\left(1-q\left(x_{n}\right)\right)^{b}}{\sum_{x_{j}\in D}\left(1-q\left(x_{j}\right)\right)^{b}}\quad\forall x_{n}\in D.$ (4)
Finally, $kD$ samples are drawn independently with replacement from $D$ using weights $w$, where $k$ was set to $1.5$ to allow for upsampling of the training data. To avoid oversampling specific samples we set the maximum number of selections to $10$.
6.2 Results
As explained in Section 4 we evaluate our method using the decoding strategy that led to the largest model collapse in each model, pure sampling and temperature sampling for GPT-2 and SmolLM2 respectively. In addition, we also implement the experiments with top-$k$ decoding to allow for direct comparison.

### v2

@@ 4 The impact of decoding strategies on model collapse
We carry out recursive training as described in Section 3.2 on the open-ended text generation task by fine-tuning LLMs on the WikiText-2 dataset Merity et al. (2016) with GPT-2 Radford et al. (2019) and SmolLM2 Allal et al. (2025). The Wikipedia articles are segmented into non-overlapping chunks of $512$ tokens, where the first $256$ are used as the context ($\mathbf{x}$), and the remaining $256$ as the continuation ($\hat{\mathbf{x}}$). We conduct full fine-tuning for $1$ epoch and, to avoid information leakage between generation and training, define cross-entropy loss only on the generated sequence of each sample Dohmatob et al. (2024b). Additional details can be found in Appendix B.3. As described in section 3.2, we recursively train language models on the previous generation’s output. We evaluate a range of decoding strategies to assess the effect on model collapse: greedy decoding ($\tau=0$), $5$-way beam search, pure sampling ($\tau=1$), temperature ($\tau=0.9$), top-$k$ (with $k=50$), and nucleus sampling ($\eta=0.95$). The hyperparameter settings for these decoding strategies were based on recommendations from prior work Holtzman et al. (2020); Shumailov et al. (2024); Garces Arias et al. (2025).
Figure 1 depicts the perplexity and evaluation accuracy on the WikiText-2 test set for every model generation. Additionally, we obtain scores for the qualitative metrics using the outputs generated by the model (i.e. $\{\hat{\mathbf{x}}\}_{s=1}^{n}$ of $\mathcal{D}_{\text{S}}^{i}$ in Algorithm 1), and enumerate them in Table 1 for generations 00 and $9$. We first notice that deterministic decoding results in the most significant model collapse. While stochastic sampling methods appear to follow a linear relationship in model degradation across generations, the rate of collapse is accelerated with greedy decoding and beam search which then starts to plateau in later generations. At generation 00, deterministic strategies yield significantly less fluent and more repetitive text, with MAUVE scores less than $5\%$ and diversity scores less than $20\%$. The gap in generation quality between deterministic and stochastic strategies in the open-ended text generation task has been explored in related literature Holtzman et al. (2020); Pillutla et al. (2021). Here we see how it leads to exacerbated model collapse. While deterministic methods are rarely used in open-ended generation, we included them for completeness and due to the choice of beam-search for the experiments of Shumailov et al. (2024), but exclude them in subsequent experiments due to their unrealistic collapse.
@@
Pure sampling approaches appear to have affected the language models in contrasting ways. For GPT-2, sampling directly from the probability distribution leads to diverse and fluent text in generation 00, but training recursively on these outputs yields the worst test perplexity among all stochastic methods, and generated text that is significantly dissimilar to human outputs (MAUVE of $7.18$). With SmolLM2, on the other hand, pure sampling led to the smallest drop in model task performance and the closest similarity to the human reference on average across all metrics.
@@
Importantly, stochastic decoding with temperature leads to the most repetitive outputs after recursive training. We observed a ${}70\%$ decrease in diversity for both models and the farthest distribution semantically from the human reference for SmolLM2, showing significant model collapse. Performance with top-$k$ sampling, on the other hand, was consistent across models, with the smallest drop in diversity and Self-BLEU, the closest resulting text to the human reference, and a smaller drop in test performance compared to nucleus sampling.
@@
In our subsequent experiments on preventing model collapse, we seek to validate that our method can work in the most extreme scenario while also facilitating direct comparison between models. For this reason, we evaluate the models using the worst-performing stochastic decoding method (pure sampling for GPT-2 and temperature sampling for SmolLM2) and top-$k$ decoding due to its consistent performance across models.
5 Preventing model collapse
So far we have carried out recursive training in a setting where models are trained exclusively on the outputs of the previous generation without implicitly including any human-generated samples. We now turn our focus to the partially synthetic setting, a more realistic scenario where human data make up a portion of the training dataset and the synthetic data is a mix of the samples produced across generations. The training dataset for generation $i$, $\mathcal{D}^{i}$, consists of the aggregation of $3$ samples:
$\begin{split}\mathcal{D}^{i}\sim&\,\,{\text{sample}_{i\geq 1}}\,\,\mathcal{D}_{\text{H}},\alpha\\ &\,\,{\text{sample}_{i\geq 1}}\,\,\mathcal{D}_{\text{S}}^{i},\beta\\ &\,\,{\text{sample}_{i\geq 2}}\,\,\left\{\mathcal{D}_{\text{S}}^{i-1},\dots,\mathcal{D}_{\text{S}}^{1}\right\},\frac{\gamma}{(i-1)}\,,\end{split}$ (3)
where $\alpha$, $\beta$, and $\gamma\!\in\![0,1]$ are mixing coefficients that affect the distribution of human and machine-generated data as well as the proportion of cross-generational data in the training set.
We explore the following settings: (i) fully synthetic $\left(\alpha\!=\!0,\beta\!=\!1,\gamma\!=\!0\right)$, where training data consists entirely of synthetic samples from the previous generation, (ii) partially synthetic $\left(\alpha\!>\!0,\beta\!=\!1,\gamma\!=\!0\right)$, where the same proportion of human data is added to the training data at every generation, and (iii) partially synthetic with cross-generational data $\left(\alpha\!>\!0,\beta\!>\!0,\gamma\!>\!0\right)$ which is the same as above, but with the synthetic data accumulated across generations. We evaluate our method in the partially synthetic setting and vary the mixing coefficients $\alpha$, $\beta$, and $\gamma$, however, our method does not assume access to the values of the mixing coefficients and hence the data distribution. To prevent model collapse when the origin of each training sample is unknown, we train a machine-generated detector that estimates the likelihood of text origin (section 5.1). We then use this information to conduct weighted sampling (section 5.2) that ultimately mitigates model collapse.
GPT-2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SmolLM2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Model collapse mitigation with GPT-2 or SmolLM2 under partially synthetic recursive training $\left(\alpha\!=\!1,\beta\!=\!1,\gamma\!=\!0\right)$ for generations 00 to $9$. The baseline is equivalent to training on all the data in the pool and the Oracle performance represents a perfect machine-generated text detector that filters all synthetic samples.
@@ 5 Preventing model collapse
RoBERTa .$982$940.940.940.940 .$940$$846$$806$804.804.804.804
DeBERTav3 .$971$$\mathbf{.954}$$\mathbf{.954}$$817$$812$810.810.810.810
ModernBERT .$\mathbf{.986}$948.948.948.948 .$948$$\mathbf{.943}$$\mathbf{.861}$$\mathbf{.860}$
Table 2: Machine-generated text detection performance. Accuracy (Acc.) and F1-score are macro-averages.
5.1 Machine-generated text detection performance
We trained and evaluated RoBERTa, DeBERTav3 and ModernBERT on the MAGE dataset Li et al. (2024), a machine-generated text detection benchmark based on documents from $10$ domains which have been used to generate text from $27$ LLMs. We adopt the preset training / validation / test splits ($80\%$/$10\%$/$10\%$). We also test on the more demanding out-of-distribution test set that contains human-curated text from $4$ unseen domains and machine-generated samples by an unseen LLM (GPT-4). Each model was fine-tuned for 5 epochs using a binary cross-entropy loss with early stopping. More details on the model training can be found in Appendix A. Performance is enumerated in Table 2. ModernBERT yielded the best classification performance on both aforementioned test sets with an AUC of .986.986.986.986 and .943.943.943.943, respectively. This is comparable to the top-performing model evaluated by Li et al. (2024), Longformer, which achieved an in-distribution and out-of-distribution AUC of .99.99.99.99 and .94.94.94.94, respectively.
5.2 Informed sampling of training data
Our goal is to sample a new training dataset at each iteration of recursive training which does not result in model collapse. We consider that a language model has collapsed if the addition of synthetic samples in the training data leads to degraded model performance compared to training exclusively on the human samples. Given text inputs $\mathbf{x}_{i}$, $i\!\in\!\{1,\dots,n\}$, with an unknown distribution of human and synthetic samples, the machine-generated text detector yields predictions $q\!\left(\mathbf{x}_{i}\right)$, $\forall i$.
Our approach deploys weighted sampling with replacement, using the detector’s prediction probabilities as weights. As the detector has been trained on an unbalanced dataset ($29\%$ human samples), the predictions are biased towards attributing text as machine-generated, reflected in the optimal classification threshold of $0.8674$ (00/$1$: human/machine-generated). To ameliorate this, we apply a bias term $b\!\geq\!1$ (see Appendix C.1) to the probabilities, followed by normalising the weights using
$w_{i}=\frac{\left(1-q\!\left(\mathbf{x}_{i}\right)\right)^{b}}{\sum_{j=1}^{n}\left(1-q\!\left(\mathbf{x}_{j}\right)\right)^{b}}\,,$ (4)
where $w_{i}\!\in\![0,1]$ denotes the weight for sample $\mathbf{x}_{i}$, $\forall i$. From the $n$ weighted training samples, we draw $k\!\times\!n$ samples with replacement. We have set $k$ to $1.5$ to allow for a $50\%$ upsampling of the training data. To avoid oversampling of specific samples we also set the maximum number of selections for an individual sample to $10$. In this way, we obtain a revised set of samples that we use in our recursive training regime.
@@ 5.2 Informed sampling of training data
GPT-2 top-$k$5.5.5.5, $1$, 00 ${\color[rgb]{0,0,1}-7.28\%}$ ${\color[rgb]{0,0,1}+2.37\%}$ ${\color[rgb]{0,0,1}+4.54\%}$ ${\color[rgb]{0,0,1}-1.77\%}$ ${\color[rgb]{0,0,1}+0.46\%}$ ${\color[rgb]{0,0,1}+9.03\%}$
.5.5.5.5, .5.5.5.5, .5.5.5.5 ${\color[rgb]{0,0,1}-5.94\%}$ ${\color[rgb]{0,0,1}+1.72\%}$ ${\color[rgb]{0,0,1}+2.43\%}$ ${\color[rgb]{0,0,1}-4.28\%}$ ${\color[rgb]{0,0,1}+3.99\%}$ ${\color[rgb]{0,0,1}+10.16\%}$
$1$, $1$, 00 ${\color[rgb]{0,0,1}-4.45\%}$ ${\color[rgb]{0,0,1}+1.49\%}$ ${\color[rgb]{0,0,1}+3.59\%}$ ${\color[rgb]{0,0,1}-3.58\%}$ ${\color[rgb]{0,0,1}+1.36\%}$ ${\color[rgb]{0,0,1}+6.76\%}$
sampling .5.5.5.5, $1$, 00 ${\color[rgb]{0,0,1}-7.41\%}$ ${\color[rgb]{0,0,1}+1.50\%}$ ${\color[rgb]{1,0,0}-1.30\%}$ ${\color[rgb]{1,0,0}+36.71\%}$ ${\color[rgb]{0,0,1}+74.06\%}$ ${\color[rgb]{0,0,1}+50.23\%}$
.5.5.5.5, .5.5.5.5, .5.5.5.5 ${\color[rgb]{0,0,1}-6.54\%}$ ${\color[rgb]{0,0,1}+1.50\%}$ ${\color[rgb]{1,0,0}-1.07\%}$ ${\color[rgb]{1,0,0}+21.39\%}$ ${\color[rgb]{0,0,1}+16.38\%}$ ${\color[rgb]{0,0,1}+25.45\%}$
$1$, $1$, 00 ${\color[rgb]{0,0,1}-25.65\%}$ ${\color[rgb]{0,0,1}+0.96\%}$ ${\color[rgb]{1,0,0}-0.71\%}$ ${\color[rgb]{1,0,0}+20.70\%}$ ${\color[rgb]{0,0,1}+16.42\%}$ ${\color[rgb]{0,0,1}+20.99\%}$
SmolLM2 top-$k$5.5.5.5, $1$, 00 ${\color[rgb]{0,0,1}-3.56\%}$ ${\color[rgb]{0,0,1}+0.40\%}$ ${\color[rgb]{0,0,1}+10.51\%}$ ${\color[rgb]{0,0,1}-5.42\%}$ ${\color[rgb]{0,0,1}+2.29\%}$ ${\color[rgb]{1,0,0}-2.50\%}$
.5.5.5.5, .5.5.5.5, .5.5.5.5 ${\color[rgb]{0,0,1}-3.50\%}$ ${\color[rgb]{0,0,1}+0.53\%}$ ${\color[rgb]{0,0,1}+7.37\%}$ ${\color[rgb]{0,0,1}-5.38\%}$ ${\color[rgb]{1,0,0}-0.71\%}$ ${\color[rgb]{1,0,0}-2.30\%}$
$1$, $1$, 00 ${\color[rgb]{0,0,1}-2.52\%}$ ${\color[rgb]{0,0,1}+0.48\%}$ ${\color[rgb]{0,0,1}+6.07\%}$ ${\color[rgb]{0,0,1}-4.77\%}$ ${\color[rgb]{0,0,1}+6.03\%}$ ${\color[rgb]{1,0,0}-1.97\%}$
temperature .5.5.5.5, $1$, 00 ${\color[rgb]{0,0,1}-3.06\%}$ ${\color[rgb]{0,0,1}+0.19\%}$ ${\color[rgb]{0,0,1}+19.77\%}$ ${\color[rgb]{0,0,1}-6.86\%}$ ${\color[rgb]{1,0,0}-13.09\%}$ ${\color[rgb]{1,0,0}-2.76\%}$
.5.5.5.5, .5.5.5.5, .5.5.5.5 ${\color[rgb]{0,0,1}-2.37\%}$ ${\color[rgb]{0,0,1}+0.27\%}$ ${\color[rgb]{0,0,1}+11.79\%}$ ${\color[rgb]{0,0,1}-6.43\%}$ ${\color[rgb]{0,0,1}+13.72\%}$ ${\color[rgb]{1,0,0}-0.98\%}$
$1$, $1$, 00 ${\color[rgb]{0,0,1}-1.49\%}$ ${\color[rgb]{0,0,1}+0.27\%}$ ${\color[rgb]{0,0,1}+11.16\%}$ ${\color[rgb]{0,0,1}-7.11\%}$ ${\color[rgb]{1,0,0}-10.49\%}$ ${\color[rgb]{1,0,0}-2.64\%}$
@@
Table 3: Percentage of change in data quality when using our proposed mitigation strategy versus the baseline. Results are shown for top-$k$ decoding and sampling/temperature for different values of $\alpha$, $\beta$, and $\gamma$ (blue / red: positive / negative results, $\uparrow$ / $\downarrow$: higher / lower is better).
5.3 Results on collapse prevention
As explained in Section 4, we assess our approach by adopting the decoding strategy that caused the most significant model collapse, i.e. pure sampling for GPT-2 and temperature sampling for SmolLM2, and for a more direct comparison, we also conduct experiments using top-$k$ decoding. At each generation $i$, we compare against the baseline of training on all samples in the pool of data $\mathcal{D}^{i}$ (Eq. 3). We also provide an “Oracle” performance which represents a perfect machine-generated text detector that filters all synthetic samples.